solve this equation using the square root property
please help me by showing how to work this out.
3x^2 = 8

Answer by KnightOwlTutor(293)

3x^2 = 8
First divide both sides by 3
x^2=8/3
Take the square root of both numbers
x= sq root of 8 =2(sq rt 2)/sq rt of 3 note that 8=4*2 the square root of 4 is 2
Need to rationalize the denomiator multiply the numerator and denomiator by sq rt of 3
2sq rt of 6/3